OpenStudy (anonymous):

Find the other endpoint of a line segment with one endpoint at (-3, 5) and whose midpoint is (5, 2).

OpenStudy (anonymous):

\[(\frac{ (y1+y2) }{ 2 },\frac{ (x1+x2) }{ 2 })=(5,2)\] also you have x1=-3,y1=5 , can you solve now ?

OpenStudy (anonymous):

yes! thank you!

OpenStudy (anonymous):

you are welcome :)